# Mapple: A Domain-Specific Language for Mapping Distributed Heterogeneous Parallel Programs

### v1

@@ Abstract.
Optimizing parallel programs for distributed heterogeneous systems remains a complex task, often requiring significant code modifications. Task-based programming systems improve modularity by separating performance decisions from core application logic, but their mapping interfaces are often too low-level. In this work, we introduce Mapple, a high-level, declarative programming interface for mapping distributed applications. Mapple provides transformation primitives to resolve dimensionality mismatches between iteration and processor spaces, including a key primitive, decompose, that helps minimize communication volume. We implement Mapple on top of the Legion runtime by translating Mapple mappers into its low-level C++ interface. Across nine applications, including six matrix multiplication algorithms and three scientific computing workloads, Mapple reduces mapper code size by 14$\times$ and enables performance improvements of up to 1.34$\times$ over expert-written C++ mappers. In addition, the decompose primitive achieves up to 1.83$\times$ improvement over existing dimensionality-resolution heuristics. These results demonstrate that Mapple simplifies the development of high-performance mappers for distributed applications.
@@ 1. Introduction
Optimizing the performance of parallel programs on modern distributed heterogeneous systems (where nodes combine different types of processors, such as CPUs and GPUs) remains a complex and labor-intensive task despite decades of research efforts. Programmers must navigate intricate low-level programming interfaces and often make substantial code modifications to achieve high performance. As a result, even experts in high-performance computing find performance tuning a persistent challenge.
@@
Two dominant programming paradigms exist today. The first is the Message Passing Interface (MPI)+X model (Gropp et al., 1999; Sanders and Kandrot, 2010; Chandra, 2001; Dagum and Menon, 1998; Yang et al., 2011) where MPI is used for distributed-memory parallelism (across computers), and X is typically a shared-memory programming model used for intra-node parallelism (within one computer), such as OpenMP or CUDA. This model gives programmers fine-grained control over performance-critical aspects but significantly compromises productivity. Achieving performance gains often requires significant changes throughout the application code, making it difficult to optimize without inadvertently affecting application logic. The second paradigm is task-based programming systems (Chamberlain et al., 2007; Augonnet et al., 2009; Fatahalian et al., 2006; Kale and Krishnan, 1993; Duran et al., 2011; Kaiser et al., 2014; Slaughter et al., 2015), which decouple performance decisions from the core application logic. These systems provide a dedicated interface for specifying performance-related policies, such as task and data placement across machines, enabling developers to tune program performance independently of application logic. This separation of concerns enhances modularity and makes performance tuning more systematic and less error-prone.
In task-based systems, the programming interface responsible for performance-critical decisions is known as the mapping interface, and the concrete implementation of these decisions is called a mapper. A central aspect of mapping is to partition tasks and data and assign them to processors in a distributed system. Distributed algorithms—such as those for matrix multiplication (Cannon, 1969; Van De Geijn and Watts, 1997; Choi et al., 1994; Solomonik and Demmel, 2011; Agarwal et al., 1995; Kwasniewski et al., 2019)—rely on carefully crafted tensor partitioning and processor mappings. For instance, the Cannon’s algorithm (Cannon, 1969) can experience up to a 3.5$\times$ slowdown when runtime heuristics are used instead of the intended mapping strategy.
@@
A major limitation of existing mapping interface designs is their low-level nature and lack of abstraction, which exposes users to the complexity and intricacies of the underlying runtime system. These interfaces are often tightly coupled with the runtime, making them difficult to use, especially for application developers unfamiliar with their internals. For instance, defining a new mapper in the Chapel framework (Chamberlain et al., 2007) requires implementing 19 functions according to 33 pages of documentation (Chamberlain et al., 2010, 2011).
In this work, we introduce Mapple, a new mapping interface design. Mapple provides a high-level abstraction that hides low-level runtime details while still exposing performance-critical mapping decisions. Compared to the original interface, Mapple reduces code size by 14$\times$ without sacrificing performance (see Section 6.1).
@@
The core challenge in any mapping interface targeting distributed applications is to determine how a multi-dimensional iteration space, defined by the algorithm’s logical loop structure, maps onto a multi-dimensional processor space, defined by the machine’s hierarchical architecture (e.g., a cluster of $n$ nodes with $m$ GPUs per node). This mapping process is directly responsible for the amount of data movement across processors, which is a key factor in achieving high performance on distributed systems. This challenge is compounded by the fact that these two spaces often differ in dimensionality, making the task of writing an effective mapping inherently complex.
To address the potential mismatch in dimensionality between the two spaces, Mapple introduces a set of transformation primitives that operate on the processor space. Building on a theoretical analysis, we identify a key primitive, decompose, which assists users in writing mappers that reduce data communication. Using this primitive, Mapple achieves up to 83% performance improvement over a commonly used heuristic for resolving dimensionality differences.
While Mapple’s transformation primitives on the processor space may appear similar to classical loop transformations, such as tiling, fusion, and fission (Wolf, 1992; Chen et al., 2008; Ragan-Kelley et al., 2013; Chen et al., 2018a), they differ fundamentally in both purpose and scope. Traditional loop transformations aim to improve locality and parallelism within a single shared-memory system. In contrast, Mapple focuses on mapping iterations across a distributed processor space. Likewise, existing scheduling languages’ approach of separating schedules from algorithms may share some similarities to Mapple’s design (Zhang et al., 2018; Senanayake et al., 2020; Ahrens et al., 2022). However, they do not address the problem of mapping iteration spaces onto processor spaces with potentially mismatched dimensionality, nor do they explicitly minimize data movement across heterogeneous processors.
We implement Mapple on top of the Legion (Bauer et al., 2012) parallel programming system. As a high-level interface, Mapple mappers are translated into Legion’s low-level C++ mapping interface. To evaluate Mapple, we apply it to a diverse set of nine applications, including six advanced distributed matrix multiplication algorithms and three scientific computing workloads. Mapple reduces the lines of code required to implement these mappers by 14$\times$ compared to hand-written C++ mappers, without introducing any observable performance overhead. In addition, we demonstrate that Mapple enables effective performance tuning: mappers written in Mapple outperform expert-crafted C++ mappers by up to 1.34$\times$.
@@
    A high-level programming interface for mapping distributed applications to heterogeneous machines.
@@
    A set of transformation primitives for resolving dimensionality mismatches between iteration and processor spaces, including a key primitive, decompose, which helps minimize data communication volume.
@@
    An implementation of Mapple that reduces mapper code size by 14$\times$ compared to the low-level interface, without incurring additional overhead.
@@
    A comprehensive evaluation showing that mappers written in Mapple can outperform expert-written C++ mappers by up to 1.34$\times$, and that the decompose primitive yields up to 1.83$\times$ performance improvement.
@@ 2. Design Goals and Non-Goals
IndexTaskMap loop0 block2d
Region task_init arg0 GPU FBMEM
Layout task_finish arg1 CPU C_order
GarbageCollect systolic arg2
@@
void map(const Task& task,
@@
Goal: A High-Level Interface Supporting Explicit Mapping and Optimization
We aim to provide a high-level programming interface that abstracts away the complexities of the underlying low-level runtime. This design allows users to focus on the logic of mapping decisions rather than the intricate details of system internals. These mapping decisions generally include:
  • how to partition and distribute loops across distributed machines;
  • where to place data in memory;
  • how to lay out data in memory;
  • whether to perform garbage collection; and
  • how to make scheduling decisions.
@@
To support this, we introduce explicit, modular abstractions and transformation primitives that make such decisions directly programmable and easy to change. For example, block2d specifies loop distribution, Region and Layout control data placement and layout, GarbageCollect handles memory management, and Backpressure guides scheduling. As shown in Figure 1, this approach achieves the same functionality as the C++ implementation but with significantly less code and greater clarity. We comprehensively quantify the reduction of code sizes in Section 6.1.
Moreover, our design enables systematic exploration of different mapping strategies. Programmers can easily experiment with different mappers, often outperforming hand-tuned C++ mappers. Figure 1(a) presents only a subset of Mapple’s features. We present the full features of Mapple in Section 7.1 and demonstrate its impact on performance tuning in Section 6.2.
Non-Goal: Portability across systems
@@
We do not aim to support seamless portability across different systems. This non-goal reflects the inherent differences among distributed runtimes, which often prevent direct reuse of implementations. Nonetheless, we believe the ideas and abstractions introduced here can inform the design of other distributed systems.
3. The Core Challenge: Mapping Iteration Space to Processor Space
@@
A central challenge for any mapping interface targeting distributed applications is efficiently assigning computation to hardware. In these settings, the algorithm defines a multi-dimensional iteration space, representing the logical structure of loops and computations. In contrast, the hardware exposes a multi-dimensional processor space, reflecting its hierarchical architecture—for example, a cluster of $n$ nodes with $m$ GPUs per node. The core problem, known as index mapping, is to map the iteration space onto the processor space in a way that minimizes data movement, which is critical for achieving high performance on distributed systems. Notably, the iteration and processor spaces often differ not only in size but also in dimensionality, adding further complexity to the mapping task.
@@
In light of this challenge, we first survey existing approaches to index mapping. Motivated by their limitations, we introduce Mapple’s transformation primitives designed for index mapping in distributed systems. We defer the discussion of how Mapple handles dimensionality mismatches between iteration and processor spaces, using the decompose primitive, to Section 4.
@@ 3.1. Existing Approaches
Existing task-based systems support mapper development through one of three design styles, as illustrated in Figure 2. The enumeration-based interface (Adaptive MPI, 2025) allows users to explicitly specify how each tile of the distributed array is mapped to a processor. While this interface is expressive, it lacks generality: enumeration-based mappers are hard-coded to a specific size of inputs or machines.
In contrast, keyword-based designs (Keyword Distribution, 2025) provide a higher-level abstraction by letting users choose from a predefined set of standard distributions (e.g., BlockDist for block distribution). However, this approach sacrifices flexibility—the fixed keyword-based options cannot express non-standard mapping strategies. For example, none of the distributions used in matrix-multiplication algorithms in Section 6.1 can be represented using keyword-based interfaces.
The programmatic approach provides users with a low-level programming interface composed of classes and functions that must be implemented to control mapping behavior. While this design offers flexibility, it exposes significant system detail and complexity. For instance, defining a new mapping in Chapel requires implementing 19 functions based on 33 pages of documentation (Chamberlain et al., 2010, 2011), and Legion’s programmatic interface directly exposes the runtime’s internal abstractions, accompanied by 19 pages of documentation (Bauer et al., 2012). In our experience, only users with deep expertise in the underlying runtime system are able to successfully develop mappers using such interfaces.
Figure 2. Three existing interface designs for mapping iteration space to processor space: the enumeration-based, keyword-based, and programmatic approaches.
@@ 3.2. Motivating Examples
We present three example mappers in Mapple to illustrate how iteration spaces are mapped onto processor spaces. First, we describe a standard block distribution in Section 3.2.1 to introduce the core concepts. Next, we discuss a custom cyclic distribution in Section 3.2.2, which is not supported in keyword-based mapping interfaces. Finally, we use a mapper from Solomonik’s 2.5D matrix multiplication algorithm (Solomonik and Demmel, 2011) in Section 3.2.3 to demonstrate the complexity introduced by mismatched dimensionality between the iteration and processor spaces.
@@ 3.2.1. Standard Block Distribution
Block distribution is a common default strategy widely supported by parallel programming systems. We use it here to illustrate the core concepts of index mapping in Mapple. In Mapple, users define mappers by writing a function that maps each iteration point in the iteration space to a target processor in the processor space. Figure 3 shows the function block2D, which implements a block distribution for an iteration space of size $(6,6)$ (that is, $60\leq x<$ 6 and $60\leq y<$ 6) and a processor space of size $(2,2)$, representing two nodes with two GPUs per node. The function assigns each iteration point to a specific processor.
The GPU processor space can be accessed using Machine(GPU). In the mapper function, the iteration point ipoint, iteration space ispace, and processor space size m.size are all 2D integer tuples. These values are used to compute the processor indices node_idx and gpu_idx, which specify the destination processor for each iteration point. As shown in Figure 3, the shaded iteration point $(2,3)$ is mapped to node 0 and GPU 1. The function returns a specific processor for each iteration point, resulting in a block distribution over the processor space.
Figure 3. Block mapping from the iteration space of $(6,6)$ to the processor space of $(2,2)$ indicating a machine with 2 nodes and 2 GPUs per node. A node index and a GPU index within the node indicate a specific GPU processor. The shaded iteration point $(2,3)$ is mapped to node 0 and GPU 1.
@@ 3.2.2. Custom Cyclic Distribution
We showcase a custom distribution expressible in Mapple that is not yet supported by most other parallel programming systems, which typically restrict users to a fixed set of standard, keyword-based distributions. This non-standard cyclic distribution is inspired by a simplified variant of a distributed matrix multiplication algorithm. Figure 4 illustrates the mapping function linearCyclic. In the resulting distribution, the subdiagonal iteration points (shaded in the iteration space) are mapped to the first processor (shaded in the processor space).
The mapper begins with the merge transformation primitive, which fuses dimension 0 and dimension 1 of the original 2D processor space into a single dimension. We will formally define the merge primitive in Section 3.3. After applying this transformation, the resulting processor space m becomes one-dimensional with size 4. In the mapping function, the 2D iteration point is first linearized, then assigned to processors using a round-robin distribution over the 4 processors.
Figure 4. A custom cyclic distribution. The merge primitive transforms the 2D processor space into a 1D space. The mapping function linearizes each 2D iteration point and applies a round-robin distribution over the resulting 1D processor space.
@@ 3.2.3. Mapping Under Dimensionality Mismatch
Figure 5. A mapper illustrating the complexity caused by mismatched dimensionality between the iteration and processor spaces, as required by the Solomonik’s algorithm on a 2-node machine with 4 GPUs per node. The original 2D processor space is transformed into a 6D space using the split primitive, visualized as two 3D spaces.
We use this example to illustrate the complexity introduced by a mismatch between the dimensionality of the iteration space and the processor space, which in turn raises deeper research questions. Figure 5 shows a mapper for the Solomonik’s algorithm executed on a 2-node machine, where each node has 4 GPUs. The iteration space is three-dimensional, and the algorithm specifies a hierarchical distribution: the iteration space is first partitioned along the x-axis so that each node handles half of the x-dimension, and within each node, the 4 GPUs perform a 2D block distribution over the y-z plane. This example highlights how mismatched dimensions require careful coordination between hierarchical and multi-dimensional mapping strategies.
There is a mismatch in dimensionality between the iteration space, which is three-dimensional, and the initial processor space, which is two-dimensional. To enable the mapping required by the algorithm, we apply the split transformation primitive four times (highlighted in red in the code). The first two splits align the node-level processor dimension with the x-axis of the iteration space, while the last two splits align the GPU-level processor dimensions with the y and z axes. This results in a six-dimensional processor space, which we visualize as two separate 3D spaces: one for nodes of size $(2,1,1)$ and one for GPUs of size $(1,2,2)$. We formally define the split transformation primitive in Section 3.3.
This example also raises deeper research questions. Can we automatically determine the appropriate splitting factors? What underlying principles guide the mapping decisions in matrix multiplication algorithms? Is it possible to address dimensionality mismatches through a generalized transformation? We explore these questions in Section 4, where we introduce a new primitive called decompose to provide a unified solution.

### v2

# Mapple: A Domain-Specific Language for Mapping Distributed Programs

@@ Abstract.
[…]
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Optimizing parallel programs on modern distributed systems remains complex and labor-intensive despite decades of research. Programmers must navigate intricate low-level programming interfaces and often make substantial code modifications to achieve high performance, making tuning challenging even for experts.
@@
Two dominant programming paradigms exist today. The first is the Message Passing Interface (MPI)+X model (gropp1999using; sanders2010cuda; chandra2001parallel; dagum1998openmp; yang2011hybrid) where MPI is used for distributed-memory parallelism (across computers), and X is typically a shared-memory programming model for intra-node parallelism (within one computer), such as OpenMP or CUDA. This model gives programmers fine-grained control over performance-critical aspects but significantly compromises productivity. Achieving performance gains often requires substantial changes throughout the application code, making it difficult to optimize without inadvertently affecting application logic. The second paradigm is task-based programming systems (chamberlain2007parallelchapel; augonnet2009starpu; fatahalian2006sequoia; kale1993charm++; duran2011ompss; kaiser2014hpx; slaughter2015regent), which separate performance decisions from core application logic. Tasks are functions whose dependencies are analyzed by the runtime and scheduled across processors. These systems offer interfaces for specifying performance policies such as task and data placement, allowing developers to tune performance independently of application logic. This separation improves modularity and makes tuning more systematic and less error-prone.
[…]
@@
A major limitation of existing mapping interface designs is their low-level nature and lack of abstraction, which exposes users to the complexity and intricacies of the underlying runtime system. These interfaces are often tightly coupled with the runtime, making them difficult to use, especially for application developers unfamiliar with runtime internals. For instance, defining a new mapper in the Chapel framework (chamberlain2007parallelchapel) requires implementing 19 functions according to 33 pages of documentation (chamberlain2010user; chamberlain2011authoring); the Legion mapping interface has more user-implementable callbacks (bauer2012legion).
[…]
@@
A core challenge in any mapping interface targeting distributed applications is to determine how a multi-dimensional task space, defined by the algorithm’s set of tasks to execute, maps onto a multi-dimensional processor space, defined by the machine’s hierarchical architecture (e.g., a cluster of $n$ nodes with $m$ GPUs per node). This mapping process is directly responsible for the amount of data movement across processors, which is a key factor in achieving high performance on distributed systems. This challenge is compounded by the fact that these two spaces often differ in dimensionality, making writing an effective mapping inherently complex.
[…]
While inspired by array programming systems such as APL (iverson1962programming), ZPL (lin1993zpl), and NumPy (harris2020array), Mapple extends these ideas to distributed task-based programs. The processor space transformation is central to the decompose primitive, which has no counterpart in prior work. Moreover, Mapple provides a richer set of mapping decisions, including processor selection, memory placement, data layout, garbage collection, and load balancing (see Section 7). These capabilities enable Mapple to support a broader range of distributed applications, including tensor algebra and scientific simulations, going beyond prior distributed scheduling systems such as Distributed Halide (denniston2016distributed), which are restricted to stencil computations and CPU-only environments.
While Mapple’s transformation primitives resemble classical loop transformations such as tiling, fusion, and fission (wolf1992improving; chen2008framework; ragan2013halide; chen2018tvm), they differ in purpose and scope. Loop transformations schedule loops, whereas Mapple schedules sets of tasks across distributed machines. Likewise, Mapple shares the goal of separating performance decisions from algorithms with prior scheduling languages (zhang2018graphit; senanayake2020sparse; ahrens2022autoscheduling), but these do not handle dimensionality mismatches between index spaces or explicitly minimize inter-processor data movement.
[…]
@@
    A high-level programming interface for mapping distributed applications.
@@
    A set of transformation primitives for resolving dimensionality mismatches between task and processor spaces, including a key primitive, decompose, which helps minimize data communication volume.
@@
[…]
@@
[…]
@@ 2. Design Goals and Non-Goals
IndexTaskMap tiles block2d
Region task_init * GPU FBMEM
Layout task_finish * CPU C_order
GarbageCollect systolic *
@@
void slice(const Task& task,
@@
Our goal is to provide a high-level programming interface that abstracts away the complexities of the underlying runtime, allowing users to focus on mapping logic rather than system internals. Specifically, Mapple enables users to place tasks across distributed machines, determine where data resides in memory, control data layout, manage garbage collection, and make scheduling decisions. We introduce abstractions and transformation primitives that make such decisions directly programmable and easy to change. For example, block2d specifies task space distribution, Region and Layout control data placement and layout, GarbageCollect handles memory management, and Backpressure guides scheduling. The Mapple mapper in Figure 1 achieves the same functionality as the C++ implementation from prior work (yadav2022distal) but with significantly less code and greater clarity. We quantify the reduction in code size in Section 6.1.
Figure 1(a) presents only a subset of Mapple’s features. We present the complete feature set of Mapple in Section 7. The effectiveness of Mapple for performance tuning is demonstrated in Section 6.2. We provide a full Mapple mapper in the supplementary material (Section A.1).
@@
We do not aim to support portability across different runtime systems. This non-goal reflects the inherent differences among distributed runtimes, which often prevent direct reuse of implementations. Nonetheless, we believe the ideas and abstractions introduced here can inform the design of mapping interfaces for other distributed systems.
3. A Core Challenge: Mapping Task Spaces to Processor Spaces
@@
A key challenge for any mapping interface in distributed applications is assigning tasks to processors. The algorithm defines a multi-dimensional task space consisting of indexed tasks, while the hardware defines a multi-dimensional processor space, capturing its hierarchical structure (e.g., a cluster of $n$ nodes with $m$ GPUs per node). The core problem, index mapping, is to define a function mapping task indices in the task space to processor indices in the processor space that minimizes data movement, which is crucial for high performance on distributed systems. Differences in dimensionality between the two spaces further complicate this task.
Task-based systems vary in how they define task spaces. Some do it implicitly (e.g., through iteration spaces of loops) and some do it explicitly (e.g., Legion’s index tasks). We assume we are given one or more task spaces to map.
@@
We first survey existing approaches to index mapping. Motivated by their limitations, we introduce Mapple’s transformation primitives designed for index mapping in distributed systems. We defer the discussion of how Mapple handles dimensionality mismatches between task and processor spaces, using the decompose primitive, to Section 4.
Figure 2. Three existing interface designs for mapping task space to processor space: the enumeration-based, keyword-based, and programmatic approaches.
@@ 3.1. Existing Approaches
Existing task-based systems support mapper development through one of three design styles, as illustrated in Figure 2. The enumeration-based interface (ampi) allows users to explicitly specify how each tile of the distributed array is mapped to a processor. While this interface is expressive, it lacks generality: enumeration-based mappers are hard-coded to a specific size of inputs or machines.
In contrast, keyword-based designs (keyword) provide a higher-level abstraction by letting users choose from a predefined set of standard distributions (e.g., BlockDist for block distribution). However, this approach sacrifices flexibility—the fixed keyword-based options cannot express non-standard mapping strategies. For example, none of the distributions used in the matrix-multiplication algorithms in Section 6.1 can be represented using keyword-based interfaces.
The programmatic approach defines an interface of classes and methods that must be implemented to control mapping behavior. While this design offers flexibility, existing programmatic interfaces expose significant system detail and complexity. For instance, defining a new mapping in Chapel requires implementing 19 functions based on 33 pages of documentation (chamberlain2010user; chamberlain2011authoring), and Legion’s programmatic interface directly exposes the runtime’s internal abstractions, accompanied by 19 pages of documentation (bauer2012legion). In our experience, only users with substantial expertise in the underlying runtime system are able to successfully develop mappers using such interfaces.
@@ 3.2. Motivating Examples
Figure 3. Block mapping from the task space $(6,6)$ to the processor space $(2,2)$, a machine with 2 nodes and 2 GPUs per node. A node index and a GPU index within the node name a specific GPU processor. The shaded index point $(2,3)$ is mapped to node 0 and GPU 1.
We present three example mappers in Mapple to illustrate how task spaces are mapped onto processor spaces. The task space comes from the application code, which is separate from the mapper. We show how tasks and task spaces are defined in Section A.1 of the supplementary material.
Below, we describe a standard block distribution in Section 3.2.1 to introduce the core concepts. Next, we discuss a custom cyclic distribution in Section 3.2.2, which is not supported in keyword-based mapping interfaces. Finally, we use a mapper from Solomonik’s 2.5D matrix multiplication algorithm (solomonik2011communication) in Section 3.2.3 to demonstrate the complexity introduced by mismatched dimensionality between the task and processor spaces.
@@ 3.2.1. Standard Block Distribution
Block distribution is a common default strategy widely supported by parallel programming systems. We use it here to illustrate the core concepts of index mapping in Mapple. In Mapple, users define mappers by writing a function that maps each point in the task space to a target processor in the processor space. Figure 3 shows the function block2D, which implements a block distribution for a task space of size $(6,6)$ (that is, $0\leq x<6$ and $0\leq y<6$) and a processor space of size $(2,2)$, representing two nodes with two GPUs per node. The function assigns each point to a specific processor.
In Mapple, the GPU processor space can be accessed by Machine(GPU), which is a 2D tuple whose shape, representing the number of nodes and GPUs per node, is automatically inferred by the runtime. Consequently, users do not need to specify these dimensions explicitly, allowing Mapple mappers to remain portable across machines of different sizes. In the mapper function, the index point ipoint, task space ispace, and processor space size m.size are all 2D integer tuples. These values are used to compute the processor indices node_idx and gpu_idx, which specify the destination processor for each index point. As shown in Figure 3, the shaded point $(2,3)$ is mapped to node 0 and GPU 1. The function returns a specific processor for each index point, resulting in a block distribution over the processor space.
Figure 4. A custom cyclic distribution. The merge primitive transforms the 2D processor space into a 1D space. The mapping function linearizes each 2D index point and applies a round-robin distribution over the resulting 1D processor space.
@@ 3.2.2. Custom Cyclic Distribution
We present a custom distribution expressible in Mapple but unsupported by most parallel programming systems that limit users to predefined keyword-based schemes. This non-standard cyclic distribution, inspired by a simplified variant of a matrix multiplication algorithm, is implemented by the mapping function linearCyclic in Figure 4, where the subdiagonal index points (shaded) map to the first processor.
The mapper first applies the merge primitive, which fuses dimensions 0 and 1 of the 2D processor space into a single dimension (formally defined in Section 3.3). The resulting processor space m is one-dimensional with size 4. The 2D task index points are then linearized and assigned to processors in a round-robin fashion.
@@ 3.2.3. Mapping Under Dimensionality Mismatch
Figure 5. A mapper illustrating the dimensionality mismatch between task and processor spaces in the Solomonik’s algorithm on a 2-node, 4-GPU-per-node machine. The original 2D processor space is transformed to 6D via the split primitive, shown as two 3D spaces.
We give an example illustrating the complexity introduced by a mismatch between the dimensionality of the task space and the processor space. Figure 5 shows a mapper for the Solomonik’s algorithm executed on a 2-node machine, where each node has 4 GPUs. The task space is three-dimensional, and the algorithm specifies a hierarchical distribution: the task space is first partitioned along the x-axis so that each node handles half of the x-dimension, and within each node, the 4 GPUs perform a 2D block distribution over the y-z plane.
The task space is three-dimensional, while the processor space is initially two-dimensional. To support the mapping required by the algorithm, we apply the split primitive four times (marked in red in the code). The first two splits align the node dimension with the task space’s x-axis, and the last two align the GPU dimensions with the y- and z-axes. This produces a six-dimensional processor space, visualized as two 3D spaces: one for nodes $(2,1,1)$ and one for GPUs $(1,2,2)$. The split primitive is formally defined in Section 3.3.
This example also raises deeper research questions. Can we automatically determine the appropriate splitting factors? What underlying principles guide the mapping decisions when resolving mismatched dimensions? Is it possible to address dimensionality mismatches through a generalized transformation? We explore these questions in Section 4, where we introduce a new primitive called decompose to provide a unified solution.